OpenStudy (anonymous):

Simplified form of 4/3y-2 - 5/9y^2-4? Thanks in advance to helpers c:

OpenStudy (anonymous):

@scorpions786 ? :)

OpenStudy (anonymous):

Help :'c

OpenStudy (anonymous):

\[\frac{ 4 }{ 3y-2 } - \frac{ 5 }{ (3y-2)(3y+2) }\] by multiplying the first with (3y+2)/(3y+2) \[\frac{ 4(3y+2) }{ 3y-2(3y+2) } -\frac{ 5 }{ (3y-2)(3y+2) }\] \[\frac{ 4(3y+2) -5 }{ (3y+2)(3y-2) }\] \[\frac{ 12y+8 -5 }{ (3y+2)(3y-2) }\] \[\frac{ 12y+3 }{ (3y+2)(3y-2) }\] sorry for that time I don't know if it's right or not

OpenStudy (anonymous):

That turned out perfect. Thank you! Don't worry about the time, I just wasn't sure if you were still present or not :p

OpenStudy (anonymous):

Are you a student?

OpenStudy (anonymous):

\[\frac{ 12y+3 }{ 9y^{2}-4 }\]

OpenStudy (anonymous):

I'm

OpenStudy (anonymous):

You're what?

OpenStudy (anonymous):

student

OpenStudy (anonymous):

What grade? I'm in 11th :)

OpenStudy (anonymous):

I duno what grade means in my country the system is different ... I'm in the 2nd college year ..

OpenStudy (anonymous):

Oh, alright, I understand. You are what's considered a Sophomore in college in the US. :) I'm in my 3rd year of high school. Next year is my last and then I'll be off to college. c: What is your country?

OpenStudy (anonymous):

Feel free to message me :D I'm going to close this message so I can ask others :) Catch you soon